undeadknight26 (undeadknight26):

Zen wants to fill a window box with soil. The length of the window box is 5 feet, width is 6 inches, and height is 9 inches. Each bag of soil contains 0.625 cubic foot of soil. How many bags of soil will Zen need to fill the window box completely? [1 foot = 12 inches]

undeadknight26 (undeadknight26):

@agent0smith please help me.

jigglypuff314 (jigglypuff314):

volume of box = (length) (width) (height) = 5 * 6 * 9 = 270 how many bags of 0.625 can go into 270? 270/0.625 =

undeadknight26 (undeadknight26):

the answer is 432?

jigglypuff314 (jigglypuff314):

That's what I got :)

undeadknight26 (undeadknight26):

Can u help me with 1 more? The equation below shows the perimeter f(x) of several rectangular paintings having width x inches: f(x) = 40 + 2x The perimeter of each painting is less than 92 inches. Which of the following best describes the domain of f(x)? All real numbers greater than 0 and less than 26 All real numbers greater than 0 and less than 92 All integers greater than 0 and less than 26 All integers greater than 0 and less than 92 i thought this was c...

jigglypuff314 (jigglypuff314):

0 < x < 26 part I know for sure is correct :) idk if it ever said it had to be integers?

undeadknight26 (undeadknight26):

I dunno this is the whole question.

jigglypuff314 (jigglypuff314):

well if the width can be something like half an inch, then A if not then C would be correct.

undeadknight26 (undeadknight26):

I dunno.

OpenStudy (phi):

For your posted question, you have to use the same units. a box 5 feet long and width 6 inches and height 9 inches does not mean a volume of 5*6*9 (5*6*9 is what you do if the sides were 5 ft, 6 ft and 9 ft)

jigglypuff314 (jigglypuff314):

hi phi :) we were discussing this one: The equation below shows the perimeter f(x) of several rectangular paintings having width x inches: f(x) = 40 + 2x The perimeter of each painting is less than 92 inches. Which of the following best describes the domain of f(x)?

OpenStudy (phi):

which is why they bother to mention (hint, hint) 12 inches is 1 foot.

jigglypuff314 (jigglypuff314):

oh right! I hadn't read properly >,< thanks phi!

OpenStudy (phi):

For *** The equation below shows the perimeter f(x) of several rectangular paintings having width x inches: f(x) = 40 + 2x *** the domain are the legal x value. x is the width. It seems you must have a width bigger than zero so 0 < x is a good start. they say the max perimeter ( f(x) ) is less than 92 so you know f(x) < 92 replace f(x) with 40+2x and simplify 40+2x < 92 2x < 52 x < 26 so it looks like 0 < x < 26

OpenStudy (phi):

There is no reason the width *has* to be integer. (We are not counting coins or eggs, we are measuring a distance)

undeadknight26 (undeadknight26):

so that means it is C?

jigglypuff314 (jigglypuff314):

all real numbers so A

undeadknight26 (undeadknight26):

oh okay thx :D

jigglypuff314 (jigglypuff314):

and I messed up the question before...

jigglypuff314 (jigglypuff314):

if was 6in = 0.5 ft 9in = 0.75ft so 5 * 0.5 * 0.75 = ___ / 0.625 =

undeadknight26 (undeadknight26):

ohhhhh ok...let me jot that down.